### v1

@@ 5 The Charging Scheme
The first step is to move the demands of $P^{(1)}_{L^{\prime}}$ from $X_{Y}\cap L^{\prime}$ to $X_{F}$. Specifically, each vertex $x_{y}\in X_{Y}\cap L^{\prime}$ will send $P^{(1)}_{L^{\prime}}(x_{y},k)=\sum_{i=1}^{\ell}{P_{L^{\prime}_{i}}(x_{y},k)}$ units of commodity $k$ to $X_{F}$, for every $k\in\mathcal{K}$. As mentioned above, this is at most $2\alpha$ for each vertex.
@@
We use Property 2 of Theorem 4.10 for the set $Y$, and “move” the demands that reside at these vertices to vertices in $X_{F}$. Property 2 says that there exists a flow which can be routed with congestion $O(1)$ in $G^{\prime}[S^{\prime}]$ such that every node $x_{y}\in X_{Y}$ sends $\mu_{\uptau}(x_{y})$ units of mass, which is at least $\uptau$ (if $y\in B$, then this is tight, otherwise $\mu_{\uptau}(x_{y})=1\geq\uptau$), and every node in $X_{F}$ receives at most $2$ units of mass. Every such $x_{y}\in X_{Y}$ has at most $2\alpha$ units of mass that it should send. Therefore, by scaling the given flow we can route this mass to nodes in $X_{F}$ with congestion $O(\frac{\alpha}{\uptau})$. By Fact 2.7 we get that this routing is $\Omega(\frac{\uptau}{\alpha})$-respected by $G^{\prime}[S^{\prime}]$. In particular, we can replace the demand state $P^{(1)}_{L^{\prime}}$ with the updated demand state $P^{(2)}_{L^{\prime}}$ after this routing, such that $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha})$-respects $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$. To be precise, we use the fact that $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$ is a valid demand state (recall Fact 3.19).
@@
After this step, any node in $X_{F}$ has load of at most $2\cdot\frac{2\alpha}{\uptau}+2\alpha\leq\frac{6\alpha}{\uptau}$ units in $P^{(2)}_{L^{\prime}}$ (as we scaled the flow by $\frac{2\alpha}{\uptau}$ and there was a load of at most $2\alpha$ in the beginning). Note that at this stage, $P^{(2)}_{L^{\prime}}$ is supported only on the vertices $X_{F}$ (since $P^{(1)}_{L^{\prime}}$ was supported on $X_{F}\cup X_{Y}$ and we moved the demands on $X_{Y}\cap L^{\prime}$). This is similar to Remark 3.20. Note that $P^{(2)}_{L^{\prime}}$ is not supported only on $L^{\prime}$, as some vertices of $X_{F}$ may lie outside of $L^{\prime}$.
Next, we update $P^{(2)}_{L^{\prime}}$ by an all-to-all multi-commodity flow on the set $X_{F}$, in order to spread the demand state $P^{(2)}_{L^{\prime}}$ uniformly on $X_{F}$. This is similar to Remark 3.20. In other words, consider the average demand vector $N_{F}\mathrel{\mathop{:}}=\frac{1}{|X_{F}|}\sum_{v\in X_{F}}{P^{(2)}_{L^{\prime}}(v)}$. Theorem 4.1 gives that $X_{F}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}[S^{\prime}]$. Using Remark 3.6, we have that the all-to-all flow problem on $X_{F}$, [^12] Recall from Definition 2.2 that this is the flow problem where each vertex in $X_{F}$ sends $\frac{1}{|X_{F}|}$ of its unique commodity to all other vertices in $X_{F}$. is $\Omega(\frac{1}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. If we scale this problem by $O(\frac{\alpha}{\uptau})$, we get that the updated demand state is $P^{(3)}_{L^{\prime}}$, where each demand vector is $N_{F}$. The valid demand state $P^{(2)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$ is therefore $\Omega(\frac{\uptau}{\alpha\log n})$-respected by $G^{\prime}[S^{\prime}]$. Using Fact 3.17, we get that $P^{(1)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$ is $\Omega(\frac{\uptau}{\alpha\log n})$-respected by $G^{\prime}[S^{\prime}]$.
Finally, due to Fact 3.19, $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$ is a valid demand state and therefore:
$N_{F}=\frac{1}{|X_{F}|}\sum_{v\in X_{F}}{P^{(2)}_{L^{\prime}}(v)}=\frac{1}{|X_{F}|}\sum_{v\in V^{\prime}}{P^{(2)}_{L^{\prime}}(v)}=\frac{1}{|X_{F}|}\sum_{v\in V^{\prime}}{P^{(1)}_{L^{\prime}}(v)}=\frac{1}{|X_{F}|}\sum_{v\in L^{\prime}}{P^{(1)}_{L^{\prime}}(v)}.$
@@
Denote by $B^{R}\subseteq E(S,V\setminus S)$ the set of boundary edges that are incident to a node in $R$. Note that $B^{R}$ may not be all of $B$, as some boundary edges can be incident to a node in $L$, as long as these boundary edges are also in $Y$ (see Figure 4). Let $\tilde{Y}\mathrel{\mathop{:}}=Y\setminus B^{R}$.
@@
A key observation is that because $L$ is node in the hierarchical decomposition tree $T$, and the original $P$ is respected by $T$, the amount of mass that has to cross the cut $(L,V\setminus L)$ is bounded by the capacity of this cut. This is the reason we included $L$ in the hierarchical decomposition tree. We first prove this slightly more general claim:
@@
For each commodity $k\in\mathcal{K}$, using Invariant 5.4(2), we have:
@@
where the first equality holds since $P^{(1)}$ is supported on $L^{\prime}$. Because the updates that moved the demand vectors from $X_{Y}$ to $X_{F}$ and the all-to-all routings do not change the total demand (i.e., the difference $P^{(3)}_{L^{\prime}}-P^{(1)}_{L^{\prime}}$ is a valid demand state), we also get
@@
    $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\log n})$-respects $P^{(3)}_{L^{\prime}}-P^{(4)}_{L^{\prime}}$.
@@
By Property 2 of Theorem 4.10 and since $X_{F}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}[S^{\prime}]$, the multi-commodity flow problem which “sends” one unit of flow from every $x_{y}\in X_{\tilde{Y}}\subseteq X_{Y}$ to a uniform distribution over $X_{F}$ is $\Omega(\frac{\uptau}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. By reversing these flows, we can distribute the demands of $P^{(3)}_{L^{\prime}}$ remaining on $X_{F}$ uniformly on the nodes in $X_{\tilde{Y}}$. By Corollary 5.11, each node $x_{e}\in X_{\tilde{Y}}$ receives at most one unit of demand. The new demand state $P^{(4)}_{L^{\prime}}$ satisfies that $P^{(3)}_{L^{\prime}}-P^{(4)}_{L^{\prime}}$ is $\Omega(\frac{\uptau}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. ∎
@@
Finally, by Fact 3.17, we can see that $G^{\prime}[S^{\prime}]$ $\gamma$-respects $P^{(1)}_{S^{\prime}}-P^{(5)}_{S^{\prime}}$, where
@@
In the context of oblivious routing schemes, the construction of [racke2014computing] was oblivious in the sense that the flow paths they construct between each pair of vertices $s,t$ were independent of the demands.
Similarly, in our case, the charges on the edges of $(B,W)$ may be defined for each pair $s,t$, independently of the demands of $P$. The construction guarantees that for any valid demand state $P$, which is $\Omega(1)$-respected by $G$ (and thus by $T$), if we add up the “oblivious” charges incurred by each $s,t$ pair, scaled up by the demands of $P$, we will incur a charge of $O(\log^{3}n)$ on each edge.
@@ 6 Improved Tree Cut-Sparsifier
    $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha\log n})$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$.
@@
We amend this using a variation on the “refinement phase” technique of [bienkowski2003practical, harrelson2003polynomial, RackeS14],131313[harrelson2003polynomial] called it the “bandwidth phase” while [bienkowski2003practical] called it the “assure-precondition” subroutine. to ensure that the all-to-all flow problem (recall Definition 2.2) between the boundary edges of the sub-clusters we recur to, is respected by the sub-clusters, meaning that the boundary edges expand in these sub-clusters (see Remark 3.6). [^14] This requirement is weaker than what was ensured in [bienkowski2003practical, harrelson2003polynomial, RackeS14]. First, we only require the all-to-all flow problem to be respected by the sub-cluster, not routable as in [bienkowski2003practical, harrelson2003polynomial]. Additionally, they ensured that a slightly stronger flow problem is routable, which implies the all-to-all problem on the boundary edges is routable (but not conversely). However, the boundary edges being expanding will be enough for our purposes.
@@
The resulting scheme is a two-level decomposition: given an input cluster $S$, we first run one level of the merge phase decomposition described in Theorem 6.1, with $\uptau=1$. This partitions $S$ to sub-clusters $\{S_{1},\ldots,S_{k}\}$, such that $|S_{i}|\leq\frac{2}{3}|S|$. This merge-phase partition of $S$ corresponds to $2$ layers in the hierarchical decomposition (see Figure 5). On each $S_{i}$ we then run the refinement phase described below. This partitions $S_{i}$ into $\{S_{i,1},\ldots,S_{i,k_{i}}\}$, such that the boundary edges of $S_{i,j}$ (i.e., $E(S_{i,j},V\setminus S_{i,j})$) expand within $G^{\prime}[S_{i,j}^{\prime}]$. Formally, we denote the boundary edges of $S_{i,j}$ as $B_{i,j}\mathrel{\mathop{:}}=E(S_{i,j},V\setminus S_{i,j})$ and prove that $X_{B_{i,j}}$ $\Omega\left(\frac{1}{f(S_{i,j})\cdot\log n}\right)$-expands in $G^{\prime}[S_{i,j}^{\prime}]$, where $f$ is defined in Theorem 6.2. The reason we need the merge phase is that we have no guarantee that the size of $S_{i,j}$ is smaller than $S_{i}$ by a constant factor. Therefore, we repeat by performing an alternation between the merge phase on each $S_{i,j}$ and then the refinement phase on each produced sub-cluster, etc.
@@
Even though $\uptau=1$, in each sub-cluster $S_{i,j}$ created by the refinement phase, we can send the demands to a uniform distribution on the boundary edges of $S_{i,j}$, so positive and negative demands cancel out and we are left with at most a load of $1$ on each boundary edge. This effectively “resets” the load on each boundary edge so it does not accumulate when we go up along the hierarchical tree. This scheme allows us to improve the charge incurred, in total over all merge phase sub-clusters, from $O(\log^{3}n)$ (by setting $\uptau=\frac{1}{\log n}$) to $O(\log^{2}n)$ ($\uptau=1$) per cut edge. As for the refinement phase sub-clusters, we will show that they incur a total of $O(\log^{2}n\log\log n)$ charge per cut edge.
@@
The main difference from [bienkowski2003practical, harrelson2003polynomial, RackeS14] is that we are able to run the refinement phase in near-linear time. Additionally, we avoid dealing with signaled events from recursive sub-calls, leading to a simpler algorithm. Explicitly, we prove the following theorem:
@@
    With high probability, for each $i$, the sub-cluster $G^{\prime}[S_{i}^{\prime}]$, $\Omega\left(\frac{1}{f(S_{i})\cdot\log n}\right)$-respects the all-to-all flow problem [^15] See Definition $2.$ Additionally, recall the definition of a graph respecting a flow problem in Definition 2.6. on $X_{B_{i}}$. Here $f(S_{i})\mathrel{\mathop{:}}=c_{f}\log\left(\frac{\sigma}{|S_{i}|}\right)\cdot\log\log n$, where $c_{f}\geq 1$ is a constant which will be defined below.
@@
    We can route in $G^{\prime}[S^{\prime}]$ a multi-commodity flow problem where each inter-cluster split node $x_{f}\in X_{F}$ (where $F=\{(u,v)\in E\mid u\in S_{i},v\in S_{j},i\neq j\}$) sends one unit of flow, and each boundary split node in $X_{B}$ (where $B=E(S,V\setminus S)$) receives at most $O(1)$ units of flow. This can be routed with congestion $O(\log n)$.
@@
Our construction begins with the merge phase (Theorem 6.1) on $S=V$ with $\uptau=1$. Then, we run the refinement phase (Theorem 6.2) on each of the resulting sub-clusters $S_{i}$, with $\sigma=|S|$. We proceed with the merge phase with $\uptau=1$ on each resulting sub-cluster $S_{i,j}$, and then with the refinement phase on $S_{i,j,k}$, etc. In each subsequent run of the refinement phase, we choose $\sigma$ to be the size of the previous cluster created by a refinement phase, i.e. the parent of the current sub-cluster (e.g., when running on $S_{i,j,k}$, we take $\sigma=|S_{i,j}|$, not $|S_{i,j,k}|$.). Because $S_{i}$ is a sub-cluster created in a merge phase, it will satisfy $|S_{i}|\leq\frac{2}{3}\cdot|V|$, ensuring the required condition on $\sigma$. We proceed recursively in this manner, running both stages on successively smaller clusters $S$, stopping when $|S|=1$. Let $T$ be the resulting hierarchical decomposition. Note that because the merge phase reduces the size of the sub-clusters by a constant fraction, the depth of $T$ is logarithmic in $n$.

### v2

@@ 5 The Charging Scheme
The first step is to move the demands of $P^{(1)}_{L^{\prime}}$ from $X_{Y}\cap L^{\prime}$ to $X_{F}$. Specifically, each vertex $x_{y}\in X_{Y}\cap L^{\prime}$ will send $|P^{(1)}_{L^{\prime}}(x_{y},k)|=|\sum_{i=1}^{\ell}{P_{L^{\prime}_{i}}(x_{y},k)}|$ units (these units may be negative) of commodity $k$ to $X_{F}$, for every $k\in\mathcal{K}$. As mentioned above, this is at most $2\alpha$ for each vertex.
@@
We use Property 2 of Theorem 4.9 for the set $Y$, and “move” the demands that reside at these vertices to vertices of $X_{F}$. Property 2 says that there exists a flow which can be routed with congestion $O(1)$ in $G^{\prime}[S^{\prime}]$ such that every node $x_{y}\in X_{Y}$ sends $\mu_{\uptau}(x_{y})$ units of mass, which is at least $\uptau$ (if $y\in B$, then this is tight, otherwise $\mu_{\uptau}(x_{y})=1\geq\uptau$), and every node in $X_{F}$ receives at most $2$ units of mass. Every such $x_{y}\in X_{Y}$ has at most $2\alpha$ units of mass that it should send. Therefore, by scaling the given flow we can route this mass to nodes in $X_{F}$ with congestion $O(\frac{\alpha}{\uptau})$. By Fact 2.7 we get that this routing is $\Omega(\frac{\uptau}{\alpha})$-respected by $G^{\prime}[S^{\prime}]$. In particular, we can replace the demand state $P^{(1)}_{L^{\prime}}$ with the updated demand state $P^{(2)}_{L^{\prime}}$ after this routing, such that $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha})$-respects $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$. To be precise, we use the fact that $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$ is a valid demand state (recall Fact 3.19).
@@
After this step, any node in $X_{F}$ has load of at most $2\cdot\frac{2\alpha}{\uptau}+2\alpha\leq\frac{6\alpha}{\uptau}$ units in $P^{(2)}_{L^{\prime}}$ (as we scaled the flow by $\frac{2\alpha}{\uptau}$ and there was a load of at most $2\alpha$ at the beginning). By Lemma 3.20(2), $P^{(2)}_{L^{\prime}}$ is supported only on the vertices $X_{F}$ (since $P^{(1)}_{L^{\prime}}$ was supported on $X_{F}\cup X_{Y}$ and we moved the demands on $X_{Y}\cap L^{\prime}$). Note that $P^{(2)}_{L^{\prime}}$ is not supported only on $L^{\prime}$, as some vertices of $X_{F}$ may lie outside of $L^{\prime}$. [^15] This can occur when there exists $e=(u,v)\in F\cap Y$ with $u,v\in R$.
Next, we update $P^{(2)}_{L^{\prime}}$ by uniformly spreading the demand vectors of $P^{(2)}_{L^{\prime}}$ on $X_{F}$. For this we use the scaled all-to-all flow problem of Lemma 3.20(1), which is defined with respect to the current demand state: each $u\in X_{F}$ sends $\frac{\left\|P^{(2)}_{L^{\prime}}(u)\right\|_{1}}{|X_{F}|}$ units of its own commodity to every $v\in X_{F}$. Let $Q_{F}$ be the corresponding demand matrix and let $N_{F}\mathrel{\mathop{:}}=\frac{1}{|X_{F}|}\sum_{v\in X_{F}}{P^{(2)}_{L^{\prime}}(v)}$ be the average demand vector that Lemma 3.20(1) yields, that is, $(P^{(2)}_{L^{\prime}})^{\uparrow Q_{F}}(v)=N_{F}$, for all $v\in X_{F}$. Let $P^{(3)}_{L^{\prime}}\mathrel{\mathop{:}}=(P^{(2)}_{L^{\prime}})^{\uparrow Q_{F}}$.
Let $\tilde{Q}_{F}$ be the all-to-all demand matrix on $X_{F}$ as in Definition 2.2. That is, $\tilde{Q}_{F}(u,v)=\frac{1}{|X_{F}|}$ for $u,v\in X_{F}$ and 0 otherwise. By Theorem 4.1, the set $X_{F}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}[S^{\prime}]$. Thus, by Remark 3.6, $\tilde{Q}_{F}$ is $\Omega(\frac{1}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. Since the rows of $Q_{F}$ differ from the rows of $\tilde{Q}_{F}$ by a factor of $O(\frac{\alpha}{\uptau})$, it follows, using Fact 3.19(2), that the valid demand state $P^{(2)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$ is $\Omega(\frac{\uptau}{\alpha\log n})$-respected by $G^{\prime}[S^{\prime}]$. Using Fact 3.16, we get that $P^{(1)}_{L^{\prime}}-P^{(3)}_{L^{\prime}}$ is also $\Omega(\frac{\uptau}{\alpha\log n})$-respected by $G^{\prime}[S^{\prime}]$.
Finally, due to Fact 3.19(1), $P^{(1)}_{L^{\prime}}-P^{(2)}_{L^{\prime}}$ is a valid demand state and therefore:
$N_{F}=\frac{1}{|X_{F}|}\sum_{v\in X_{F}}{P^{(2)}_{L^{\prime}}(v)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{=}}\frac{1}{|X_{F}|}\sum_{v\in V^{\prime}}{P^{(2)}_{L^{\prime}}(v)}=\frac{1}{|X_{F}|}\sum_{v\in V^{\prime}}{P^{(1)}_{L^{\prime}}(v)}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{=}}\frac{1}{|X_{F}|}\sum_{v\in L^{\prime}}{P^{(1)}_{L^{\prime}}(v)},$
where Equality $(2)$ holds since $P^{(2)}_{L^{\prime}}$ is supported on $X_{F}$, and Equality $(4)$ holds since $P^{(1)}_{L^{\prime}}$ is supported on $L^{\prime}$. ∎
@@
Denote by $B^{R}\subseteq E(S,V\setminus S)$ the set of boundary edges that are incident to a node in $R$. Note that $B^{R}$ may not be all of $B$, as some boundary edges can be incident to a node in $L$, as long as these boundary edges are also in $Y$. Let $\tilde{Y}\mathrel{\mathop{:}}=Y\setminus B^{R}$ (see Figure 4).
@@
A key observation is that since $L$ is node in the hierarchical decomposition tree $T$, and the original demand state $P$ is respected by $T$, the amount of mass that has to cross the cut $(L,V\setminus L)$ is bounded by the capacity of this cut. This is the reason we included $L$ in the hierarchical decomposition tree. We formalize this observation in the following claim:
@@
For each commodity $k\in\mathcal{K}$ and $i\in\{1,\ldots,\ell\}$, using Invariant 5.4(2), we have:
@@
where the first equality holds since $P^{(1)}$ is supported on $L^{\prime}$. Since the updates that moved the demand vectors from $X_{Y}$ to $X_{F}$ and the all-to-all routing do not change the total demand (i.e., the difference $P^{(3)}_{L^{\prime}}-P^{(1)}_{L^{\prime}}$ is a valid demand state), we also get for each commodity $k\in\mathcal{K}$ that
@@
    $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\log n})$-respects $P^{(3)}_{L^{\prime}}-P^{(4)}_{L^{\prime}}$. Moreover, $P^{(3)}_{L^{\prime}}-P^{(4)}_{L^{\prime}}$ is a valid demand state.
@@
By Property 2 of Theorem 4.9 and since $X_{F}$ $\Omega(\frac{1}{\log n})$-expands in $G^{\prime}[S^{\prime}]$, the demand matrix which “sends” one unit of flow from every $x_{y}\in X_{\tilde{Y}}\subseteq X_{Y}$ to a uniform distribution over $X_{F}$ is $\Omega(\frac{\uptau}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. Indeed, as done in Lemma 5.9, we can first route the demands from $X_{\tilde{Y}}$ to $X_{F}$ and then spread the demands uniformly. By reversing these demands movements, we can distribute the demands of $P^{(3)}_{L^{\prime}}$ remaining on $X_{F}$ uniformly on the nodes in $X_{\tilde{Y}}$. By Corollary 5.11, each node $x_{e}\in X_{\tilde{Y}}$ receives at most one unit of demand. The new demand state $P^{(4)}_{L^{\prime}}$ satisfies that $P^{(3)}_{L^{\prime}}-P^{(4)}_{L^{\prime}}$ is $\Omega(\frac{\uptau}{\log n})$-respected by $G^{\prime}[S^{\prime}]$. ∎
@@
Finally, by Fact 3.16, we can see that $G^{\prime}[S^{\prime}]$ $\gamma$-respects $P^{(1)}_{S^{\prime}}-P^{(5)}_{S^{\prime}}$, where
@@
In the context of oblivious routing schemes, the construction of [27] was oblivious in the sense that the flow paths they construct between each pair of vertices $s,t$ were independent of the demands.
Similarly, in our case, the charges on the edges of $(U,W)$ may be defined for each pair $s,t$, independently of the demands of $P$. The construction guarantees that for any valid demand state $P$, which is $\Omega(1)$-respected by $G$ (and thus by $T$), if we add up the “oblivious” charges incurred by each $s,t$ pair, scaled up by the demands of $P$, we will incur a charge of $O(\log^{3}n)$ on each edge.
@@ 6 Improved Tree Cut-Sparsifier
    $G^{\prime}[S^{\prime}]$ $\Omega(\frac{\uptau}{\alpha\log n})$-respects $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$. Moreover, $P^{\mathrm{before}}_{S^{\prime}}-P^{\mathrm{after}}_{S^{\prime}}$ is a valid demand state.
@@
We amend this using a variation on the “refinement phase” technique of [7, 15, 28],161616[15] called it the “bandwidth phase” while [7] called it the “assure-precondition” subroutine. to ensure that the all-to-all flow problem (recall Definition 2.2) between the boundary edges of the sub-clusters we recur to, is respected by the sub-clusters, meaning that the boundary edges expand in these sub-clusters (see Remark 3.6). [^17] This requirement is weaker than what was ensured in [7, 15, 28]. First, we only require the all-to-all flow problem to be respected by the sub-cluster, not routable as in [7, 15]. Additionally, they ensured that a slightly stronger flow problem is routable, which implies the all-to-all problem on the boundary edges is routable (but not conversely). However, the boundary edges being expanding will be enough for our purposes.
@@
The resulting scheme is a two-level decomposition: given an input cluster $S$, we first run one level of the merge phase decomposition described in Theorem 6.1, with $\uptau=1$. This partitions $S$ to sub-clusters $\{S_{1},\ldots,S_{k}\}$, such that $|S_{i}|\leq\frac{2}{3}|S|$. This merge phase partition of $S$ corresponds to $2$ layers in the hierarchical decomposition (see Figure 5). On each $S_{i}$ we then run the refinement phase described below. This partitions $S_{i}$ into $\{S_{i,1},\ldots,S_{i,k_{i}}\}$, such that the boundary edges of $S_{i,j}$ (i.e., $E(S_{i,j},V\setminus S_{i,j})$) expand within $G^{\prime}[S_{i,j}^{\prime}]$. Formally, we denote the boundary edges of $S_{i,j}$ as $B_{i,j}\mathrel{\mathop{:}}=E(S_{i,j},V\setminus S_{i,j})$ and prove that $X_{B_{i,j}}$ $\Omega\left(\frac{1}{f(S_{i,j})\cdot\log n}\right)$-expands in $G^{\prime}[S_{i,j}^{\prime}]$, where $f$ is defined in Theorem 6.2. The reason we need the merge phase is that we have no guarantee that the size of $S_{i,j}$ is smaller than $S_{i}$ by a constant factor. Therefore, we repeat by performing an alternation between the merge phase on each $S_{i,j}$ and then the refinement phase on each produced sub-cluster, etc.
@@
Even though $\uptau=1$, in each sub-cluster $S_{i,j}$ created by the refinement phase, we can send the demands to a uniform distribution on the boundary edges of $S_{i,j}$, so positive and negative demands cancel out and we are left with at most a load of $1$ on each boundary edge. This effectively “resets” the load on each boundary edge so it does not accumulate when we go up along the hierarchical tree. This scheme allows us to improve the charge incurred, in total over all merge phase sub-clusters, from $O(\log^{3}n)$ (by setting $\uptau=\frac{1}{\log n}$) to $O(\log^{2}n)$ ($\uptau=1$) per cut edge. As for the sub-clusters of the refinement phase, we show that they incur a total of $O(\log^{2}n\log\log n)$ charge per cut edge.
@@
The main difference from [7, 15, 28] is that we are able to run the refinement phase in near-linear time. Additionally, we avoid dealing with signaled events from recursive sub-calls, leading to a simpler algorithm. Explicitly, we prove the following theorem:
@@
    With high probability, for each $i$, the sub-cluster $G^{\prime}[S_{i}^{\prime}]$, $\Omega\!\left(\frac{1}{f(S_{i})\cdot\log n}\right)$-respects the all-to-all flow problem [^18] See Definition 2.2. Additionally, recall the definition of a graph respecting a flow problem in Definition 2.6. on $X_{B_{i}}$, where $B_{i}=E(S_{i},V\setminus S_{i})$. Here $f(S_{i})\mathrel{\mathop{:}}=c_{f}\log\left(\frac{\sigma}{|S_{i}|}\right)\cdot\log\log n$, where $c_{f}\geq 1$ is a constant which will be defined below.
@@
    We can route in $G^{\prime}[S^{\prime}]$ a demand matrix where each inter-cluster split node $x_{f}\in X_{F}$ (where $F=\{(u,v)\in E\mid u\in S_{i},v\in S_{j},i\neq j\}$) sends one unit of flow, and each boundary split node in $X_{B}$ (where $B=E(S,V\setminus S)$) receives at most $O(1)$ units of flow. This can be routed with congestion $O(\log n)$.
Intuition for $f$ and $\sigma$.
The quantity $f(S_{i})\mathrel{\mathop{:}}=c_{f}\log\left(\frac{\sigma}{|S_{i}|}\right)\cdot\log\log n$ plays two distinct roles in the analysis. The parameter $\sigma$ is a global scale parameter that controls how the costs incurred by the refinement phase aggregate across the levels of the hierarchical decomposition, allowing for a telescoping summation. At the same time, the quantity $f(S_{i})$ provides a local guarantee whose strength depends on the size of the sub-cluster $S_{i}$. Larger sub-clusters admit stronger all-to-all respect guarantees. This is crucial for the routing guarantee in Item (2) of Theorem 6.2, since the routing will proceed in a bottom-up manner with respect to the refinement-induced hierarchy on $S$. Intuitively, if a refinement step significantly reduces the size of a sub-cluster, then the expansion requirement can be relaxed, as the smaller cluster contributes less to congestion (as they tend to be lower at the hierarchy). Conversely, if the cluster remains large, the analysis forces a stronger expansion guarantee. Thus, the refinement phase always makes progress either by enforcing strong expansion or by substantially shrinking the cluster size.
@@
Our construction, see Algorithm 1, begins with the merge phase (Theorem 6.1) on $S=V$ with $\uptau=1$. Then, we run the refinement phase (Theorem 6.2) on each of the resulting sub-clusters $S_{i}$, with $\sigma=|S|$. We proceed with the merge phase with $\uptau=1$ on each resulting sub-cluster $S_{i,j}$ (i.e., the sub-clusters created from $S_{i}$ in the previous refinement step), and then with the refinement phase on $S_{i,j,k}$, etc. In each subsequent run of the refinement phase, we choose $\sigma$ to be the size of the previous cluster created by a refinement phase, i.e. the parent of the current sub-cluster (e.g., when running on $S_{i,j,k}$, we take $\sigma=|S_{i,j}|$, not $|S_{i,j,k}|$.). Because $S_{i}$ is a sub-cluster created in a merge phase, it will satisfy $|S_{i}|\leq\frac{2}{3}\cdot|V|$, ensuring the required condition on $\sigma$. We proceed recursively in this manner, running both stages on successively smaller clusters $S$, stopping when $|S|=1$. Let $T$ be the resulting hierarchical decomposition. Note that because the merge phase reduces the size of the sub-clusters by a constant fraction, the depth of $T$ is logarithmic in $n$.